OpenStudy (anonymous):

the point P(a, klogea) lies on the curve y =klogex where "a" and "k" are constants a) show that the equation of the tangent T to the curve at P is kx-ay +ak(logea-1)

myininaya (myininaya):

is the base e? or is that e*a?

myininaya (myininaya):

\[P(a,k*\log_e(a)) or P(a,k*\log(e*a))\]?

myininaya (myininaya):

\[y=k*\log_ex or y=\log(e*x)\] ?

OpenStudy (anonymous):

lost

myininaya (myininaya):

ok i will asumme it is base e and if it isn't correct oh well at least you will have an example

OpenStudy (anonymous):

\[y=k\ln(x)\]?

myininaya (myininaya):

\[y=k*lnx\] \[y'=k*\frac{1}{x}\] \[y'(a)=k*\frac{1}{a}\] is the slope at x=a

OpenStudy (anonymous):

so point is \[(a, k\ln(a))\] now that makes sense yes?

myininaya (myininaya):

\[y=mx+b\] \[k*lna=\frac{k}{a}*a+b\] we need to solve for b the y-intercept \[b=k*lna-k\] so the tangent line at P is \[y=\frac{k}{a}*x+k*lna-k\]

OpenStudy (anonymous):

point slope formula give \[y-k\ln(a)=\frac{k}{a}(x-a)\]

OpenStudy (anonymous):

\[y=\frac{k}{a}x-k+k\ln(a)\]

OpenStudy (anonymous):

does multiplying by a get what you want?

myininaya (myininaya):

looks like we were suppose to get slope is k

myininaya (myininaya):

so maybe we didn't interpret what she had correctly oh well

myininaya (myininaya):

i give you a medal satellite :)

OpenStudy (anonymous):

want this \[kx-ay +ak(\ln(a)-1)\]

myininaya (myininaya):

she/he whatever

myininaya (myininaya):

but where is the = sign

OpenStudy (anonymous):

oh not it is ok. just algebra

myininaya (myininaya):

wait is it y=kx-ay+ak(lna-1)

OpenStudy (anonymous):

ikr maybe it is this \[kx-ay =ak(\ln(a)-1)\]

OpenStudy (anonymous):

i bet that is it.

OpenStudy (anonymous):

multiply by a and we get it

myininaya (myininaya):

the plus sign = equal sign that would work

OpenStudy (anonymous):

yes i think so. just \[y=\frac{k}{a}x-k+k\ln(a)\] \[ay=kx-al+ak\ln(a)\] and some other stupid algebra probably does it.

OpenStudy (anonymous):

and don't make any... typos!

OpenStudy (anonymous):

@myininaya and @satellite73 THANK YOU A BILLION TIMES OVER and i deeply apologise for the confusion. like myininaya had suggested, it is base e! THANK YOU, I WOULD GIVE YOU GUYS 10000000000000000000 MEDALS EACH IF THAT WAS POSSIBLE!

myininaya (myininaya):

lol np